# More Text, Less Point: Towards 3D Data-Efficient Point-Language Understanding

### v1

@@ Introduction
Inspired by CLIP (Radford et al. 2021), multimodal large language models (MLLMs) can map different modality inputs to a text space closer to LLMs using pre-trained multimodal encoders, enabling LLMs to understand data beyond just language. Existing 3D point-language models follow a similar approach, applying LLMs to 3D understanding by learning from 3D point-text data pairs (Luo et al. 2024; Qi et al. 2024b). For example, PointLLM (Xu et al. 2023) and ShapeLLM (Qi et al. 2024a) employ pre-trained multimodal point cloud encoders (Xue et al. 2024; Qi et al. 2024a), mapping the point cloud space to the text space. This leaves the alignment of point cloud with LLMs to only align the text space with LLMs, which is relatively easier for LLMs. Finally, they propose to fine-tune the whole model with large amount of 3D-text data pairs, thus enhancing the LLMs’ 3D understanding capabilities. However, this field remains under-explored. The primary reason is that training LLMs requires billions of text data from the internet, while 3D-text pair data is scarce since 3D data itself is hard to acquire and requires expensive annotations. Consequently, the scaling law that drives LLMs success are difficult to achieve in the 3D domain, directly limiting the development of 3D foundation models.
@@
To fairly and reasonably evaluate the models, we introduce a new metric to measure the efficiency of 3D data usage, and establish a new evaluation benchmark based on open-source LLMs. Experimental results show that our GreenPLM outperforms previous models using only 12% of the 3D data. It even surpasses GPT4Point (Qi et al. 2024b) without any 3D data, maintaining extremely 3D data-efficient point-language understanding, which demonstrates the effectiveness of our approach. The contributions of this paper are as follows:
@@
To enable LLMs to understand the 3D physical world, early approaches project 3D point clouds into 2D images, relying on 2D-LLMs for comprehension (Hong et al. 2023). Point-Bind LLM (Guo et al. 2023) establishes a 3D-2D-LLM connection for better performance. However, these 2D-based methods lose crucial 3D information, leading to issues like occlusion, ambiguity, and hallucination. Recently, with the availability of large-scale 3D-text data (Luo et al. 2024; Qi et al. 2024b) and multimodal encoders, methods like PointLLM (Xu et al. 2023) and ShapeLLM (Qi et al. 2024a) connect point encoders with LLMs and fine-tune the 3D Point Cloud-LLMs (3D-LLMs) using vast amounts of 3D-text data. Unfortunately, comparing to images, 3D-text data remains extremely scarce (Objaverse-1M vs. LAION-5B) (Deitke et al. 2023; Schuhmann et al. 2022) and expensive, let alone the near infinite and free text data, making it challenging to build powerful 3D foundation models according to the scaling law. Also, training 3D-LLMs is resource-intensive, often requiring 8xA100 GPUs for hundreds of hours. Although MiniGPT-3D (Tang et al. 2024) reduces training time to 26.8h on a single GPU, the 3D data bottleneck persists. Our GreenPLM proposes to solve this 3D data bottleneck by leveraging extensive text data to compensate for the lack of 3D data, and introducing a 3-stage training strategy for effective and efficient alignment.
@@
To achieve this, we bring T3D, a 6M text dataset of 3D object descriptions and conversations. Fig. 3 shows the verb-noun distribution and a visualized word cloud. Instead of using the closed-source GPT-4 (OpenAI 2023), we use the equally powerful open-source model Qwen2-72B-Instruct (Yang et al. 2024a) to construct this dataset. We select object categories from Cap3D (Luo et al. 2024) and DiffuRank (Luo, Johnson, and Lee 2024), and we design prompts to generate 4 types of data: 1M brief descriptions, 1M detailed descriptions, 3M single-round conversations, and 1M multi-round conversations. The object descriptions help the LLMs learn rich semantic knowledge, while the conversations enable the LLMs to extract useful information from the context to improve 3D understanding. Notably, this dataset is constructed without any manual annotation or post-processing, requiring only minimal model inference cost. Each type of data follows the Caption-Question-Answer format, as shown in Tab. 1. During training, we input the Caption into the text encoder, pass the encoded tokens through a projector, and then input them along with the Question into the LLM, which outputs a response to calculate the loss against the Answer. More detailed prompts and distributions are in Appendix.
@@
For better multimodal encoder-LLM alignment and minimizing the use of 3D point-text data pairs, we propose a 3-stage training strategy, as shown in Fig. 4. Our design principle is to first use a large amount of text data to align the text encoder with the LLM via a MLP projector (Stage I and II). Then, using only a small amount of 3D point-text data pairs, we align the point cloud encoder with the LLM via the same projector (Stage III). Specifically, for each stage, we will introduce the pipeline, trainable layers, and data aspects as follows.
@@
is shown in Fig. 4(a). First, we input a text caption $\bm{D}$ of a 3D object into the pre-trained text encoder $f_{text}$, obtaining the global text embedding $\bm{C}_{t}$ as the encoder output. $\bm{C}_{t}$ is then passed through a learnable MLP projector $f_{proj}$ to connect with the LLM $f_{LLM}$. The LLM input consists of the projector output $f_{proj}(\bm{C}_{t})$, and the text tokens of an instruction prompt $\bm{I}$ such as “What is this?”. Finally, the LLM outputs a brief description $\bm{R}_{brief}$ of the 3D object, which can be used to calculate the loss with the ground-truth description. The formulas are as follows:
@@
To validate our 3D data-free capability, we compared GreenPLM-0 with the SOTA 2D-LLMs, InstructBLIP and LLaVA, as well as the 3D-2D-LLM model Point-Bind LLM (Guo et al. 2023). To evaluate GreenPLM with limited 3D data, we choose the SoTA models PointLLM (Xu et al. 2023) and MiniGPT-3D (Tang et al. 2024). For fairness, we train both using the same 90K limited 3D point-text datas.

### v2

@@ Introduction
Inspired by CLIP (Radford et al. 2021), multimodal large language models (MLLMs) can map different modality inputs to a text space closer to LLMs using pre-trained multimodal encoders, enabling LLMs to understand data beyond just language. Existing 3D point-language models follow a similar approach, applying LLMs to 3D understanding by learning from 3D point-text data pairs (Luo et al. 2024; Qi et al. 2024b). For example, PointLLM (Xu et al. 2023) and ShapeLLM (Qi et al. 2024a) employ pre-trained multimodal point cloud encoders (Xue et al. 2024; Qi et al. 2024a), mapping the point cloud space to the text space. This leaves the alignment of point cloud with LLMs to only align the text space with LLMs, which is relatively easier for LLMs. Finally, they propose to train the 3D-LLMs with large amount of 3D-text data pairs, thus enhancing the LLMs’ 3D understanding capabilities. However, this field remains under-explored. The primary reason is that training LLMs requires billions of datas, while 3D-text pair data is scarce because 3D data itself is hard to acquire and requires expensive annotations. Consequently, the scaling law that drives LLMs success are difficult to achieve in the 3D domain, directly limiting the development of 3D foundation models.
@@
To fairly and reasonably evaluate the models, we introduce a new metric to measure the efficiency of 3D data usage, and establish a new evaluation benchmark based on open-source LLMs. Experimental results show that our GreenPLM outperforms previous models using only 12% of the 3D data. It even surpasses GPT4Point (660K) (Qi et al. 2024b) without any 3D data, maintaining extremely 3D data-efficient point-language understanding, which demonstrates the effectiveness of our approach. The contributions of this paper are as follows:
@@
To enable LLMs to understand the 3D physical world, early attempt (Hong et al. 2023) projects 3D point clouds into 2D images, relying on 2D-LLMs for comprehension. However, 2D-based method lose crucial 3D information, leading to issues like occlusion, ambiguity, and hallucination. Point-Bind LLM (Guo et al. 2023) attempts to establish a 3D-2D-LLM connection, but this non-robust link leads to unstable performance. Recently, with the availability of large-scale 3D-text data (Luo et al. 2024; Qi et al. 2024b) and multimodal encoders, methods like PointLLM (Xu et al. 2023) and ShapeLLM (Qi et al. 2024a) connect point encoders with LLMs and fine-tune the 3D Point Cloud-LLMs (3D-LLMs) using vast amounts of 3D-text data. Unfortunately, compared to images, 3D-text data remains extremely scarce (LAION-5B vs. Objaverse-1M) (Schuhmann et al. 2022; Deitke et al. 2023) and expensive, let alone the near infinite and free text data, making it challenging to build powerful 3D foundation models according to the scaling law. Also, training 3D-LLMs is resource-intensive, often requiring 8xA100 GPUs for hundreds of hours. Although MiniGPT-3D (Tang et al. 2024) reduces training time to 26.8h on a single GPU, the 3D data bottleneck persists. Our GreenPLM proposes to solve this 3D data bottleneck by leveraging extensive text data to compensate for the lack of 3D data, and introducing a 3-stage training strategy for effective and efficient alignment.
@@
To achieve this, we bring T3D, a 6M text dataset of 3D object descriptions and conversations. Fig. 3 shows the verb-noun distribution and a visualized word cloud. Instead of using the closed-source GPT-4 (OpenAI 2023), we use the equally powerful open-source model Qwen2-72B-Instruct (Yang et al. 2024a) to construct this dataset. We select object categories from Cap3D (Luo et al. 2024) and DiffuRank (Luo, Johnson, and Lee 2024), and we design prompts to generate 5 types of data: 1M captions, 1M brief descriptions, 1M detailed descriptions, 3M single-round conversations, and 1M multi-round conversations. The object descriptions help the LLMs learn rich semantic knowledge, while the conversations enable the LLMs to extract useful information from the context to improve 3D understanding. Notably, this dataset is constructed without any manual annotation or post-processing, requiring only minimal model inference cost. Five types of data, totaling 6M samples in the Caption-Question-Answer format, are shown in Table 1. During training, we input the Caption into the text encoder, pass the encoded tokens through a projector, and then input them along with the Question into the LLM, which outputs a response to calculate the loss against the Answer. More detailed prompts and distributions are in Appendix.
@@
For better multimodal encoder-LLM alignment and minimizing the use of 3D point-text data pairs, we propose a 3-stage training strategy, as shown in Fig. 4. Our design principle is to first use a large amount of text data to align the text encoder with the LLM via a MLP projector (Stage I and II). Then, using only a small amount of 3D point-text datas, we align the point cloud encoder with the LLM via the same projector (Stage III). Specifically, for each stage, we will introduce the pipeline, trainable layers, and data aspects as follows.
@@
is shown in Fig. 4(a). First, we input a text caption $\bm{D}$ of a 3D object into the pre-trained text encoder $f_{text}$, obtaining the global text embedding $\bm{C}_{t}$ as the encoder output. $\bm{C}_{t}$ is then passed through a learnable MLP projector $f_{proj}$ to connect with the LLM $f_{LLM}$. The LLM input consists of the projector output $f_{proj}(\bm{C}_{t})$, and the text tokens of an instruction prompt $\bm{I}$, such as “What is this?”. Finally, the LLM outputs a brief description $\bm{R}_{brief}$ of the 3D object, which can be used to calculate the loss with the ground-truth description. The formulas are as follows:
@@
To validate our 3D data-free capability, we compared GreenPLM-0 with the SoTA 2D-LLMs, InstructBLIP and LLaVA, as well as the 3D-2D-LLM model Point-Bind LLM (Guo et al. 2023). To evaluate GreenPLM with limited 3D data, we choose the SoTA 3D-LLMs PointLLM (Xu et al. 2023) and MiniGPT-3D (Tang et al. 2024). For fairness, we train both using the same 90K limited 3D point-text datas.